# A proof of the Kim-Vu sandwich conjecture

### v1

@@ 3.1 Edge deletion process
Let $0<1/C\ll\varepsilon\ll 1$ and $d\geq C\log n$. Let $\mu$ and $C_{0}$ be such that $1/C\ll\mu\ll 1/C_{0}\ll\varepsilon$. For each $0\leq i<\binom{n}{2}-dn/2$, let
@@
Let $f\in[n]^{(2)}\setminus E(F)$. Consider $\mathbb{P}(F_{i}=F|F_{i-1}=F+f)$. If $F_{i-1}=F+f$, the probability that, for the first $j^{\prime}>\ell(i-1)$ with $e_{j^{\prime}}\in E(F_{i-1})$, the edge $e_{j^{\prime}}$ is not immediately removed from $F_{i-1}=F+f$ is, using (9) and that $x_{j^{\prime}}\sim U([0,1])$,
@@
    For each $0\leq i\leq\binom{n}{2}-dn/2-R$, and each $e\in E(F_{i-1})$,
@@
By Claim 3.2 iv) and v), B1 and B2 hold with high probability. When this happens, from Claim 3.2 i) and ii) we have that, for all $1\leq j\leq N-R$, $m(j)\leq k(j+R)\leq\ell(j+R)$. Thus, by Claim 3.2 iii), $G\subset F_{N}\subset G^{*}_{N-R}$. As $G^{*}=G^{*}_{\binom{n}{2}-M}$, when $M\geq\binom{n}{2}-N+R=dn/2+3R$, we have that $G^{*}_{N-R}\subset G^{*}$. As $M\sim\mathrm{Bin}\left((1+\varepsilon)\frac{d}{n},\binom{n}{2}\right)$ and $R=\lfloor\varepsilon dn/8\rfloor$, by a Chernoff bound (Lemma 2.6), with high probability we have $M\geq(1+\varepsilon)dn/2-R\geq dn/2+3R$. Therefore $G\subset G^{*}$ with high probability. Thus, it is left only to prove Claim 3.2 i)v) which we do in Sections 3.5.13.5.5 respectively (using Theorem 2.4).
@@
Suppose that B2 holds. Let $0\leq i\leq N-R$ and suppose, for contradiction, that $k(i+R)<m(i)$. Then, there is a set $I\subset[i+R]$ such that $|I|=R+1$ and, for each $j\in I$, $k(j)\notin\{m(i^{\prime}):i^{\prime}\in[i]\}$. Note that in particular $i\geq 1$. Take $j\in I$. As $k(j)\leq k(i+R)<m(i)$, there is some $i^{\prime}\in[i]$ such that $m(i^{\prime}-1)<k(j)<m(i^{\prime})$. From Claim 3.2 i), we have $k(i-1)\leq m(i-1)$, so therefore $j\geq i^{\prime}$.
@@
As $m(j)\leq\ell(j+R)$, there is some $j^{\prime}\leq j+R$ such that $\ell(j^{\prime}-1)<m(j)\leq\ell(j^{\prime})$. Note that either $e_{m(j)}\notin E(F_{j^{\prime}-1})$ or
@@
Note that, as $\mathcal{K}_{d}(F_{i-1})\neq\emptyset$, it follows that $\min_{f\in E(F_{i-1})}|\{K\in\mathcal{K}_{d}(F_{i-1}):f\in E(K)\}|>0$. Note also that, by double counting,
@@ 4 Main switching argument
Label vertices so that $e=u_{1}u_{2}$ and $f=v_{1}v_{2}$. Let $\mathcal{K}^{e}$ be the set of $K\in\mathcal{K}_{d}(F)$ which contain $e$ but not $f$, and let $\mathcal{K}^{f}$ be the set of $K\in\mathcal{K}_{d}(F)$ which contain $f$ but not $e$. From (19) and Claim 4.2, we have $|\mathcal{K}^{e}|>18\eta d|\{K\in\mathcal{K}_{d}(F):f\in E(K)\}|>|\mathcal{K}_{d}(F)|/n^{2}$, where have used that $\eta d\geq C_{0}$.
@@
Let $K\in\mathcal{K}^{e}\setminus\mathcal{K}^{\textrm{bad}}$. Fix $v_{1}^{\prime}\in N_{K}(v_{1})\setminus\{u_{1},u_{2},v_{2}\}$ and $v_{2}^{\prime}\in N_{K}(v_{2})\setminus\{u_{1},u_{2},v_{1},v_{1}^{\prime}\}$, for which there are at least $(d-3)(d-4)\geq(1-\eta)d^{2}$ choices, where we have used that $\eta d\geq C_{0}$. Let $v_{1}^{\prime\prime}\in N_{F\setminus K}(v_{1}^{\prime})\setminus\{u_{1},u_{2},v_{1},v_{2},v_{2}^{\prime}\}$, with at least $(1-2\eta)\delta$ choices by LABEL:prop:degreebound. Let $P_{1}$ be an $(F\setminus K,K)$-alternating $u_{1},v_{1}^{\prime\prime}$-path of length $2\ell$ which does not have any internal vertices in $\{u_{2},v_{1},v_{1}^{\prime},v_{2},v_{2}^{\prime}\}$, noting that from LABEL:prop:lowerboundKFKalternating there are at least $\left(1-2\eta\right)d^{\ell}\delta^{\ell}/n$ choices for $P_{1}$, again using that $\eta d\geq C_{0}$.
@@ 5 Upper bounds for alternating paths from a fixed vertex
We now prove the following lemma, which gives a good upper bound on the number of $(F\setminus K,K)$-alternating paths starting from a fixed vertex. (That is, unlike our ultimate goal of LABEL:prop:upperboundFKKalternating, one of the endpoints of the paths considered is not fixed). As with the next two sections, we consider the regime corresponding to $m\leq n^{1+\sigma}$ in Theorem 2.4, for some small fixed $\sigma>0$, but we prove that one of the properties, LABEL:eq:onlydegreesinFK:random in Lemma 5.3, holds more widely so that we can use it in Section 8 and also to conclude that LABEL:prop:degreebound is likely to hold. As with our results across the next three sections, instead of working in $F(n,d,m)$ here we work with the slightly easier to study random graph $F$ formed by taking $K\sim K_{d}(n)$ and adding edges independently at random with some probability $p$ chosen so that the expected number of added edges is $m$. When we apply the main results of these sections, in Section 9, at the cost of a manageable increase in the probability the events considered hold, we will easily be able to translate these results into the model $F(n,d,m)$.
@@ 5.1 Pseudorandom properties for upper bounds for paths from a vertex
To show the number of $(F\setminus K,K)$-alternating paths from a vertex are bounded we will use bounds on the degrees of $F\setminus K$ and, as discussed in Section 2.2.2, a more complicated condition on certain degree sums required for the accuracy we need. Due to this more complicated condition, we prove the lemma by induction by considering not the number of paths, but the sum of the degrees in $F\setminus K$ of the endpoints of the paths we wish to consider.
@@
For each $1\leq i\leq 10\log n$, it therefore follows from (21) that, for each $v^{\prime}\in[n]$, the number of $(K,F\setminus K)$-alternating paths with length $2i$ starting at $v^{\prime}$ is at most
$\left(1+\lambda\right)\cdot\left(1+\frac{\lambda}{\log n}\right)^{i-1}\delta^{i}d^{i}$ $\leq\left(1+2\lambda+\frac{2(i-1)\lambda}{\log n}\right)\delta^{i}d^{i}\leq\left(1+30\lambda\right)\delta^{i}d^{i},$
where we have used that $\lambda\ll 1$. Thus, LABEL:prop:upperboundFKKalternating:random:new holds. ∎
@@ 5.2 Likely properties of $F\setminus K$: degree bounds
Let $\hat{E}_{v}$ be the set of edges $e\in([n]\setminus\{v\})^{(2)}\setminus E(K)$ which contain at least one vertex in $A$. For each $xy\in\hat{E}_{v}$, let $w_{xy}=|N_{K}(x)\cap A|+|N_{K}(y)\cap A|$, so that $w_{xy}\leq 200$ as $|N_{K}(u)\cap A|<100$ for each $u\in[n]$. Note that
@@
where we have used that, for each $u^{\prime}\in[n]$, we have $d_{(K_{n}\setminus K)-v}(u^{\prime})\in\{n-2-d,n-1-d\}$. Therefore, we can take subsets $\hat{E}_{v,i}$, $i\in[200]$, in $\hat{E}_{v}$ such that $|\hat{E}_{v,i}|\geq|A|d(n-1-d)/400$ and, for each $xy\in\hat{E}_{v}$, $|\{i\in[200]:xy\in\hat{E}_{v,i}\}|=w_{xy}$. For example, if we randomly assign each $xy\in\hat{E}_{v}$ to $w_{xy}$ of the $\hat{E}_{v,i}$, $i\in[200]$, for each $xy\in\hat{E}_{v}$. Then, for each $i\in[200]$, we have $\mathbb{E}|\hat{E}_{v,i}|=\sum_{xy\in\hat{E}_{v}}\frac{w_{xy}}{200}$, and by Theorem 2.5 there is a positive probability that $|\hat{E}_{v,i}|\geq 2\mathbb{E}|\hat{E}_{v,i}|/3\geq A|d(n-1-d)/400$ for all $i\in[200]$, where we have used (23).
@@
Let $C^{\prime}$ satisfy $1/n\ll C^{\prime}\ll 1$. As $\delta=2m/n\geq\varepsilon C\log n$, we have that
@@ 6.1 Pseudorandom properties for lower bounds for path counts
Let $1/n\ll 1/C\ll 1/C_{0}\ll 1$ and $1/C\ll\sigma\ll 1$. Let $d\geq C\log n$ with $d,\delta\leq n^{\sigma}$. Let $\log^{-3}n\leq\lambda\ll 1$ satisfy $\lambda\delta\geq C_{0}\log n$. Let $K\in\mathcal{K}_{d}(n)$ and let $F$ be a graph with vertex set $[n]$, such that following properties hold.
@@
    For each $U\subset[n]$ with $|U|\leq 100\log n$, $|\{v\in[n]\setminus U:d_{F\setminus K}(v,U)\geq 10\}|\leq 100\log n$.
@@
For this, let $1\leq j\leq\ell_{1}$ and assume that (27)–(30) hold for each $1\leq i<j$.
We start by showing that, for each $1\leq a<j$,
@@
As $|X_{0}|=r$, using LABEL:eq:degreesinFK:pseud we have that $(1-\lambda)\delta r\leq\sum_{v\in X_{0}}d_{F\setminus K}(v)\leq(1+\lambda)\delta r$. Thus, by (29) applied for each $1\leq i\leq r$, we have, as $a<\ell_{1}=o(\log n)$ and $\lambda\ll 1$, that
@@
Now, for each $1\leq a<j$, using $|X_{0}|=|Y_{0}|=r$ when $a=1$ and otherwise (32) and (33) for i=a1i=a-1, by (27) with $i=a$
@@
If $j=1$, then set $X_{-1}=X^{\prime}_{0}$ and $Y_{-1}=Y^{\prime}_{0}$ (using the sets in the claim statement). Whether $j=1$, or not, we then have that $X_{2j-2}\cup Y_{2j-2}\subset N_{K}(X_{2j-3}\cup Y_{2j-3})$, and, from the properties of $X^{\prime}_{0}$ and $Y^{\prime}_{0}$ or by (31) and (34), $|X_{2j-3}\cup Y_{2j-3}|\leq 2\delta^{j-1}d^{j-2}r\leq 4|X_{2j-2}\cup Y_{2j-2}|/d$. Furthermore, either as $|X_{0}|=|Y_{0}|=r$ or by (31) with a=j1a=j-1, and as $\delta^{\ell_{1}}d^{\ell_{1}}<\lambda n/(2\cdot 10^{7})$ by the choice of $\ell_{1}$, we have that
$|X_{2j-2}\cup Y_{2j-2}|\leq 2\cdot\frac{6}{5}d^{j-1}\delta^{j-1}r<2\cdot\frac{\lambda n}{10^{7}}\cdot d^{j-1-\ell_{1}}\delta^{j-1-\ell_{1}}\leq\frac{\lambda n}{\delta d}\leq\frac{\lambda n}{100\delta\log n}.$ (35)
@@
$|E_{\geq 2,F\setminus K}(X_{2j-2}\cup Y_{2j-2})|\leq\frac{\lambda}{\log n}\cdot\delta|X_{2j-2}|\leq\frac{\lambda}{\log n}\cdot 2\delta^{j}d^{j-1}r.$ (36)
Furthermore, we have $\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})\subset\left(\cup_{a=0}^{2j-3}(X_{a}\cup Y_{a})\right)\cup N_{K}\left(\cup_{a=0}^{2j-3}(X_{a}\cup Y_{a})\right)$, and, using (31) and (34),
$|\cup_{a=0}^{2j-3}(X_{a}\cup Y_{a})|\leq 2\cdot\frac{6}{5}\sum_{a=0}^{2j-3}\delta^{\lceil a/2\rceil}d^{\lfloor a/2\rfloor}r\leq 3\delta^{j-1}d^{j-2}r\leq\frac{4}{d}|X_{2j-2}\cup Y_{2j-2}|\leq\frac{4}{d}|\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})|.$
Similarly to (35), we have $|\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})|\leq{\lambda n}/{100\delta\log n}$. Thus, from LABEL:eq:fewdoubled:F-K, we have
$e_{F\setminus K}\left(\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})\right)$ $\leq\frac{\lambda}{\log n}\cdot\delta\left|\cup_{a=0}^{2j-2}(X_{a}\cup Y_{a})\right|\leq\frac{\lambda}{\log n}\cdot 3\delta^{j}d^{j-1}r,$ (37)
@@
$|W|\leq 3r+\lambda\delta\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=1\}}\right)\cdot\delta^{j}d^{j-1}r,$ (38)
@@
$\Bigg(\sum_{v\in X_{2j-2}}d_{F\setminus K}(v)\Bigg)-|X_{2j-1}|$ $\leq|E_{\geq 2,F\setminus K}(X_{2j-2}\cup Y_{2j-2})|+e_{F\setminus K}\left(\cup_{j^{\prime}=0}^{2j-2}(X_{j^{\prime}}\cup Y_{j^{\prime}})\right)+|W|$
@@
From (39), we get the lower-bound in (27) holds for $i=j$. As $|X_{2j-1}|\leq|N_{F\setminus K}(X_{2j-2})|\leq\sum_{v\in X_{2j-2}}d_{F\setminus K}(v)$, we therefore have that (27) holds entirely for $i=j$. Furthermore, by a similar deduction to those above using LABEL:eq:degreesinFK:pseud, we have that $\frac{4}{5}d^{j-1}\delta^{j}r\leq|X_{2j-1}|\leq\frac{6}{5}d^{j-1}\delta^{j}r$. Similarly, we can deduce that
$\Bigg(\sum_{v\in Y_{2j-2}}d_{F\setminus K}(v)\Bigg)-|Y_{2j-1}|$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{j=1\}}\right)\cdot 8\cdot\sum_{v\in Y_{2j-2}}d_{F\setminus K}(v),$ (40)
that (28) holds for $i=j$, and that $\frac{4}{5}d^{j-1}\delta^{j}r\leq|Y_{2j-1}|\leq\frac{6}{5}d^{j-1}\delta^{j}r$.
Now, working similarly to above, we have that $X_{2j-1}\cup Y_{2j-1}\subset N_{F\setminus K}(X_{2j-2}\cup Y_{2j-2})$ and $|X_{2j-2}\cup Y_{2j-2}|\leq 4|X_{2j-1}\cup Y_{2j-1}|/\delta$. Furthermore, we have that
$|X_{2j-1}\cup Y_{2j-1}|\leq 2\cdot\frac{6}{5}d^{j-1}\delta^{j}r\leq\frac{2\lambda n}{10^{7}}\cdot d^{j-1-\ell_{1}}\delta^{j-\ell_{1}}\leq\left\{\begin{array}[]{ll}\frac{\lambda n}{10^{6}d\log n}&\text{ if }j<\ell_{1},\\ \frac{\lambda n}{10^{6}d}&\text{ if }j=\ell_{1}.\end{array}\right$.

### v2

@@ 3.1 Edge deletion process
Let $0<1/C\ll\varepsilon\ll 1$ and $d\geq C\log n$. Let $\mu$ and $C_{0}$ be such that $1/C\ll\mu\ll 1/C_{0}\ll\varepsilon$. Note that we can assume that $(1+\varepsilon)d/n\leq 1$, for otherwise in Theorem 1.1 we can take $p^{*}=1$. For each $0\leq i<\binom{n}{2}-dn/2$, let
@@
Let $f\in[n]^{(2)}\setminus E(F)$. Consider $\mathbb{P}(F_{i}=F|F_{i-1}=F+f)$. Observe that
$\sum_{e\in E(F+f)}|\mathcal{K}_{d}(F+f-e)|=(e(F)+1-dn/2)|\mathcal{K}_{d}(F+f)|.$ (13)
If $F_{i-1}=F+f$, the probability that, for the first $j^{\prime}>\ell(i-1)$ with $e_{j^{\prime}}\in E(F_{i-1})$, the edge $e_{j^{\prime}}$ is not immediately removed from $F_{i-1}=F+f$ is, using (9), (3.5.4) and that $x_{j^{\prime}}\sim U([0,1])$,
@@
    For each $1\leq i\leq\binom{n}{2}-dn/2-R$, and each $e\in E(F_{i-1})$,
@@
We will now show that Claim 3.1 follows from Claim 3.2. By Claim 3.2 iv) and v), B1 and B2 hold with high probability. When this happens, from Claim 3.2 i) and ii) we have that, for all $1\leq j\leq N-R$, $m(j)\leq k(j+R)\leq\ell(j+R)$. Thus, by Claim 3.2 iii), $G\subset F_{N}\subset G^{*}_{N-R}$. As $G^{*}=G^{*}_{\binom{n}{2}-M}$, when $M\geq\binom{n}{2}-N+R=dn/2+3R$, we have that $G^{*}_{N-R}\subset G^{*}$. As $M\sim\mathrm{Bin}\left((1+\varepsilon)\frac{d}{n},\binom{n}{2}\right)$ and $R=\lfloor\varepsilon dn/8\rfloor$, by a Chernoff bound (Lemma 2.6), with high probability we have $M\geq(1+\varepsilon)dn/2-R\geq dn/2+3R$. Therefore $G\subset G^{*}$ with high probability. Thus, it is left only to prove Claim 3.2 i)v) which we do in Sections 3.5.13.5.5 respectively (using Theorem 2.4).
@@
Suppose that B2 holds. Let $0\leq i\leq N-R$ and suppose, for contradiction, that $k(i+R)<m(i)$. Then, there is a set $I\subset[i+R]$ such that $|I|=R+1$ and, for each $j\in I$, $k(j)\notin\{m(i^{\prime}):i^{\prime}\in[i]\}$. Note that in particular $|I|\geq 1$. Take $j\in I$. As $k(j)\leq k(i+R)<m(i)$, there is some $i^{\prime}\in[i]$ such that $m(i^{\prime}-1)<k(j)<m(i^{\prime})$. From Claim 3.2 i), we have $k(i-1)\leq m(i-1)$, so therefore $j\geq i^{\prime}$.
@@
As $m(j)\leq\ell(j+R)$, there is some $j^{\prime}\leq j+R$ such that $\ell(j^{\prime}-1)<m(j)\leq\ell(j^{\prime})$. Note that
@@
Note also that, by double counting,
@@ 4 Main switching argument
Label vertices so that $e=u_{1}u_{2}$ and $f=v_{1}v_{2}$. Let $\mathcal{K}^{e}$ be the set of $K\in\mathcal{K}_{d}(F)$ which contain $e$ but not $f$, and let $\mathcal{K}^{f}$ be the set of $K\in\mathcal{K}_{d}(F)$ which contain $f$ but not $e$. From (20) and Claim 4.2, we have $|\mathcal{K}^{e}|>18\eta|\{K\in\mathcal{K}_{d}(F):f\in E(K)\}|>|\mathcal{K}_{d}(F)|/n^{2}$, where have used that $\eta d\geq C_{0}$.
@@
Let $K\in\mathcal{K}^{e}\setminus\mathcal{K}^{\textrm{bad}}$. Fix $v_{1}^{\prime}\in N_{K}(v_{1})\setminus\{u_{1},u_{2},v_{2}\}$ and $v_{2}^{\prime}\in N_{K}(v_{2})\setminus\{u_{1},u_{2},v_{1},v_{1}^{\prime}\}$, for which there are at least $(d-3)(d-4)\geq(1-\eta)d^{2}$ choices, where we have used that $\eta d\geq C_{0}$. Let $v_{1}^{\prime\prime}\in N_{F\setminus K}(v_{1}^{\prime})\setminus\{u_{1},u_{2},v_{1},v_{2},v_{2}^{\prime}\}$, with at least $(1-2\eta)\delta$ choices by LABEL:prop:degreebound and $\eta^{2}\delta\geq C_{0}\log{n}$. Let $P_{1}$ be an $(F\setminus K,K)$-alternating $u_{1},v_{1}^{\prime\prime}$-path of length $2\ell$ which does not have any internal vertices in $\{u_{2},v_{1},v_{1}^{\prime},v_{2},v_{2}^{\prime}\}$, noting that from LABEL:prop:lowerboundKFKalternating there are at least $\left(1-2\eta\right)d^{\ell}\delta^{\ell}/n$ choices for $P_{1}$, again using that $\eta d\geq C_{0}$.
@@ 5 Upper bounds for alternating paths from a fixed vertex
We now prove the following lemma, which gives a good upper bound on the number of $(F\setminus K,K)$-alternating paths starting from a fixed vertex. (That is, unlike our ultimate goal of LABEL:prop:upperboundFKKalternating, one of the endpoints of the paths considered is not fixed). As with the next two sections, we consider the regime corresponding to $m\leq n^{1+\sigma}$ in Theorem 2.4, for some small fixed $\sigma>0$, but we prove that one of the properties, LABEL:eq:onlydegreesinFK:random in Lemma 5.3, holds more widely so that we can use it in Section 8 and also to conclude that LABEL:prop:degreebound is likely to hold. As with our results across the next three sections, instead of working in $F(n,d,m)$ here we work with the slightly easier to study random graph $F$ formed by taking $K\sim G_{d}(n)$ and adding edges independently at random with some probability $p$ chosen so that the expected number of added edges is $m$. When we apply the main results of these sections, in Section 9, at the cost of a manageable increase in the probability the events considered hold, we will easily be able to translate these results into the model $F(n,d,m)$.
@@ 5.1 Pseudorandom properties for upper bounds for paths from a vertex
To show the number of $(F\setminus K,K)$-alternating paths from a vertex are bounded we will use bounds on the degrees of $F\setminus K$ and, as discussed in Section 2.2.2, a more complicated condition on certain degree sums required for the accuracy we need. Due to this more complicated condition, we prove the lemma by induction – considering not the number of paths, but the sum of the degrees in $F\setminus K$ of the endpoints of the paths we wish to consider.
@@
Now let $v^{\prime}$ be in $N_{F\setminus K}(v)$. Clearly if $\ell=1$ then the number of $(F\setminus K,K)$-alternating paths with length $2\ell$ starting at $v$ is $|N_{F\setminus K}(v)|\cdot d\leq(1+\lambda)\delta d$ by LABEL:eq:onlydegreesinFK. For $\ell\geq 2$, it follows from (22) that, for each $v\in[n]$, the number of $(F\setminus K,K)$-alternating paths with length $2\ell$ starting at $v$ is at most
$|N_{F\setminus K}(v)|\cdot\left(1+\lambda\right)\left(1+\frac{\lambda}{\log n}\right)^{\ell-2}\delta^{\ell-1}d^{\ell-1}\cdot d$ $\leq(1+\lambda)^{2}\left(1+\frac{2\ell\lambda}{\log n}\right)\delta^{\ell}d^{\ell}\leq\left(1+30\lambda\right)\delta^{i}d^{i},$
where we have used LABEL:eq:onlydegreesinFK and that $\ell\leq 10\log{n}$ and $\lambda\ll 1$. Thus, LABEL:prop:upperboundFKKalternating:random:new holds. ∎
@@ 5.2 Likely properties of $F\setminus K$: degree bounds
Let $\hat{E}_{v}$ be the set of edges $e\in([n]\setminus\{v\})^{(2)}\setminus E(K)$. For each $xy\in\hat{E}_{v}$, let $w_{xy}=|N_{K}(x)\cap A|+|N_{K}(y)\cap A|$, so that $w_{xy}\leq 200$ as $|N_{K}(u)\cap A|<100$ for each $u\in[n]$. Note that
@@
where we have used that, for each $u^{\prime}\in[n]$, we have $d_{(K_{n}\setminus K)-v}(u^{\prime})\in\{n-2-d,n-1-d\}$. Therefore, we can take subsets $\hat{E}_{v,i}$, $i\in[200]$, in $\hat{E}_{v}$ such that $|\hat{E}_{v,i}|\geq|A|d(n-1-d)/400$ and, for each $xy\in\hat{E}_{v}$, $|\{i\in[200]:xy\in\hat{E}_{v,i}\}|=w_{xy}$. For example, if we randomly assign each $xy\in\hat{E}_{v}$ to $w_{xy}$ of the $\hat{E}_{v,i}$, $i\in[200]$, for each $xy\in\hat{E}_{v}$. Then, for each $i\in[200]$, we have $\mathbb{E}|\hat{E}_{v,i}|=\sum_{xy\in\hat{E}_{v}}\frac{w_{xy}}{200}$, and by Theorem 2.5 there is a positive probability that $|\hat{E}_{v,i}|\geq 2\mathbb{E}|\hat{E}_{v,i}|/3\geq|A|d(n-1-d)/400$ for all $i\in[200]$, where we have used (24).
@@
Let $C^{\prime}$ satisfy $1/C\ll 1/C^{\prime}\ll 1$. As $\delta=2m/n\geq\varepsilon C\log n$, we have that
@@ 6.1 Pseudorandom properties for lower bounds for path counts
Let $1/n\ll 1/C\ll 1/C_{0}\ll 1$ and $1/C\ll\sigma\ll 1$. Let $d\geq C\log n$ with $d,\delta\leq n^{\sigma}$. Let $\log^{-3}n\leq\lambda\ll 1$ satisfy $\lambda\delta\geq C_{0}\log n$. Let $K\in\mathcal{K}_{d}(n)$ and let $F$ be a graph with vertex set $[n]$, such that $K\subset F$ and the following properties hold.
@@
    For each $U\subset[n]$ with $|U|\leq 100\log n$, $|\{v\in[n]\setminus U:|N_{F\setminus K}(v)\cap U|\geq 10\}|\leq 100\log n$.
@@
For this, let $1\leq i\leq\ell_{1}$ and assume that (28)–(31) hold for each $1\leq a<i$.
We start by showing that, for each $1\leq a<i$,
@@
As $|X_{0}|=r$, using LABEL:eq:degreesinFK:pseud we have that $(1-\lambda)\delta r\leq\sum_{v\in X_{0}}d_{F\setminus K}(v)\leq(1+\lambda)\delta r$. Thus, by (30) applied for each $1\leq a\leq i$, we have, as $i\leq\ell_{1}=o(\log n)$ and $\lambda\ll 1$, that
@@
Now, for each $1\leq a<i$, using $|X_{0}|=|Y_{0}|=r$ when $a=1$ and otherwise (33) and (34) for i=a1i=a-1, by (28) with $i=a$
@@
If $i=1$, then set $X_{-1}=X^{\prime}_{0}$ and $Y_{-1}=Y^{\prime}_{0}$ (using the sets in the claim statement). Whether $i=1$, or not, we then have that $X_{2i-2}\cup Y_{2i-2}\subset N_{K}(X_{2i-3}\cup Y_{2i-3})$, and, from the properties of $X^{\prime}_{0}$ and $Y^{\prime}_{0}$ or by (32) and (35), $|X_{2i-3}\cup Y_{2i-3}|\leq 3\delta^{i-1}d^{i-2}r\leq 4|X_{2i-2}\cup Y_{2i-2}|/d$. Furthermore, either as $|X_{0}|=|Y_{0}|=r$ or by (32) with a=i1a=i-1, and as $\delta^{\ell_{1}}d^{\ell_{1}}<\lambda n/(2\cdot 10^{7})$ by the choice of $\ell_{1}$, we have that
$|X_{2i-2}\cup Y_{2i-2}|\leq 2\cdot\frac{6}{5}d^{i-1}\delta^{i-1}r<2\cdot\frac{\lambda n}{10^{7}}\cdot d^{i-1-\ell_{1}}\delta^{i-1-\ell_{1}}\leq\frac{\lambda n}{\delta d}\leq\frac{\lambda n}{100\delta\log n}.$ (36)
@@
$|E_{\geq 2,F\setminus K}(X_{2i-2}\cup Y_{2i-2})|\leq\frac{\lambda}{\log n}\cdot\delta|X_{2i-2}|\leq\frac{\lambda}{\log n}\cdot 2\delta^{i}d^{i-1}r.$ (37)
Furthermore, we have $\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})\subset\left(\cup_{a=0}^{2i-3}(X_{a}\cup Y_{a})\right)\cup N_{K}\left(\cup_{a=0}^{2i-3}(X_{a}\cup Y_{a})\right)$, and, using (32) and (35),
$|\cup_{a=0}^{2i-3}(X_{a}\cup Y_{a})|\leq 2\cdot\frac{6}{5}\sum_{a=0}^{2i-3}\delta^{\lceil a/2\rceil}d^{\lfloor a/2\rfloor}r\leq 3\delta^{i-1}d^{i-2}r\leq\frac{4}{d}|X_{2i-2}\cup Y_{2i-2}|\leq\frac{4}{d}|\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})|.$
Similarly to (36), we have $|\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})|\leq{\lambda n}/{100\delta\log n}$. Thus, from LABEL:eq:fewdoubled:F-K, we have
$e_{F\setminus K}\left(\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})\right)$ $\leq\frac{\lambda}{\log n}\cdot\delta\left|\cup_{a=0}^{2i-2}(X_{a}\cup Y_{a})\right|\leq\frac{\lambda}{\log n}\cdot 3\delta^{i}d^{i-1}r,$ (38)
@@
$|W|\leq 3r+\lambda\delta\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=1\}}\right)\cdot\delta^{i}d^{i-1}r,$ (39)
@@
$\Bigg(\sum_{v\in X_{2i-2}}d_{F\setminus K}(v)\Bigg)-|X_{2i-1}|$ $\leq|E_{\geq 2,F\setminus K}(X_{2i-2}\cup Y_{2i-2})|+e_{F\setminus K}\left(\cup_{i^{\prime}=0}^{2i-2}(X_{i^{\prime}}\cup Y_{i^{\prime}})\right)+|W|$
@@
From (40), we get the lower-bound in (28) holds. As $|X_{2i-1}|\leq|N_{F\setminus K}(X_{2i-2})|\leq\sum_{v\in X_{2i-2}}d_{F\setminus K}(v)$, we therefore have that (28) holds entirely. Furthermore, by a similar deduction to those above using LABEL:eq:degreesinFK:pseud, we have that $\frac{4}{5}d^{i-1}\delta^{i}r\leq|X_{2i-1}|\leq\frac{6}{5}d^{i-1}\delta^{i}r$. Similarly, we can deduce that
$\Bigg(\sum_{v\in Y_{2i-2}}d_{F\setminus K}(v)\Bigg)-|Y_{2i-1}|$ $\leq\left(\frac{\lambda}{\log n}+\lambda\cdot\mathbf{1}_{\{i=1\}}\right)\cdot 8\cdot\sum_{v\in Y_{2i-2}}d_{F\setminus K}(v),$ (41)
that (29) holds, and that $\frac{4}{5}d^{i-1}\delta^{i}r\leq|Y_{2i-1}|\leq\frac{6}{5}d^{i-1}\delta^{i}r$.
Now, working similarly to above, we have that $X_{2i-1}\cup Y_{2i-1}\subset N_{F\setminus K}(X_{2i-2}\cup Y_{2i-2})$ and $|X_{2i-2}\cup Y_{2i-2}|\leq 4|X_{2i-1}\cup Y_{2i-1}|/\delta$. Furthermore, we have that
$|X_{2i-1}\cup Y_{2i-1}|\leq 2\cdot\frac{6}{5}d^{i-1}\delta^{i}r\leq\frac{2\lambda n}{10^{7}}\cdot d^{i-1-\ell_{1}}\delta^{i-\ell_{1}}\leq\left\{\begin{array}[]{ll}\frac{\lambda n}{10^{6}d\log n}&\text{ if }i<\ell_{1},\\ \frac{\lambda n}{10^{6}d}&\text{ if }i=\ell_{1}.\end{array}\right$.